OpenStudy (anonymous):

Solve for x: (3x - 1)(3x - 1) = 0

OpenStudy (anonymous):

is it just x = -1/3?

OpenStudy (anonymous):

one is the answer use foil

OpenStudy (anonymous):

the problem was originally \[9x^2 - 6x + 1 = 0\]

OpenStudy (anonymous):

i factored it

OpenStudy (anonymous):

you definitely don't want to multiply out after you factor! just write \[3x-1=0\] \[x=\frac{1}{3}\] and be done

OpenStudy (anonymous):

that is the answer